# UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

### v3

@@ Appendix C Appendix C: Experiment Detail
Count Alignment CouA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023) ITI Count Alignment evaluates how accurately the number of objects described in the text matches both the number and type of objects presented in the corresponding image. It specifically measures whether the quantity of objects mentioned in the text (e.g., two dogs, three cars, or five apples) is correctly reflected in the image, and ensures that the objects shown are of the correct type. For example, if the text mentions ’two dogs’, the image should contain two dogs, not two cats. This evaluation focuses solely on the count and type of objects, without considering their attributes or positions.
Object Alignment ObjA (Li et al. 2024a; Ghosh, Hajishirzi, and Schmidt 2023) ITI Object Alignment evaluates how accurately the objects described in a text are represented in an image. The objects mentioned in the text should be accurately presented in the image, with no omissions. For example, in the description ’ A man sitting by himself at the dinner table’, the image should feature both a man and a dinner table to achieve a high score. Missing or extra objects in the image would lower the score.
Scene Alignment ScnA (Li et al. 2024a) ITI Scene Alignment assesses the extent to which the setting or background described in a text is accurately represented in an image. It focuses on elements such as location, weather conditions, time of day, and general atmosphere rather than specific objects. For example, if a text describes ’a sunny environment with historical architecture around a town square’, the image should depict this specific setting, showing historical architecture situated within a town square context, not isolated in an open field or desert.
Attribute Alignment AttA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023) ITI Attribute Alignment assesses how well the attributes (such as color, texture, size, and shape) of objects described in the text are reflected in the image. It ensures that key features like color and size match what’s described. For instance, if the text says “a green bicycle,” the image should show a bicycle that is green. It also checks if details like texture and shape are accurate, like a “smooth, red ball” being depicted as both smooth and red. Furthermore, the relative size and proportions of objects should be consistent with the description. Overall, the image should visually match the text’s description of each object’s attributes.
Color Alignment ColA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023) ITI This criterion assesses whether the colors described in the text match the corresponding objects in the image based solely on the text description, or whether the colors of the objects conform to realistic colors. If the text specifies the color of an object, the evaluation checks whether the image reflects that color. For example, if the text says ’a blue car’, the image should show a car that matches the described color, not one of a different color.
@@
Shape Alignment ShaA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023) ITI This criterion evaluates how accurately the shapes described in the text correspond to the shapes of objects in the image, based solely on the text’s shape descriptors, or whether the shape of the objects conforms to realistic shape. The text may include common shape terms such as square, round, rectangular, oval, triangular, cubic, cylindrical, spherical, and other basic or specialized shapes like pyramidal, conical, pentagonal, teardrop, crescent, and diamond.
@@
Texture Alignment TexA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023) ITI This criterion evaluates how accurately the textures described in the text are reflected in the image, based solely on the text description, or whether the textures of the objects conform to realistic textures. It includes surface qualities such as smoothness, roughness, or granularity, as well as material types like wood, plastic, and rubber. Additionally, it may encompass tactile sensations or patterns, such as silky, polished, or other material-based descriptions. The accuracy of the texture description is assessed by comparing the textual description to the actual appearance and surface properties of the objects in the image.
@@
Advanced Alignment AdvA (Li et al. 2024a) ITI Advanced Alignment is a comprehensive evaluation aspect that assesses the ability to accurately represent complex textual descriptions in an image. It encapsulates Comparison, Negation, and Universality aspects. This means it checks whether the image correctly depicts comparisons between entities based on attributes such as number or volume; represents absence or contradiction of elements indicated by words like ’no’ or ’not’; and illustrates scenarios where every member of a group shares a specific attribute or relation.
Table 13: The definition and source of Universal Aspects. (Part 4)
Aspect Abbr Source Target Definition
@@
Comparison Cmpa (Li et al. 2024a) ITI Comparison evaluates the ability to accurately depict comparisons between entities as described in a text. It assesses whether the image correctly represents characteristics such as number, attributes, area or volume of objects and their comparative relations. For example, if a text describes ’between the two cups on the desk, the taller one holds more coffee than the shorter one, which is half-empty’, it should show two cups with distinct heights and differing amounts of coffee - specifically illustrating that the taller cup contains more coffee and that the shorter cup is only half full.
@@
Negation Nega (Li et al. 2024a) ITI Negation evaluates the ability to accurately depict the absence or contradiction of elements as described in a text. It assesses whether the image correctly represents scenarios where entities are not present, as indicated by words such as ’no’, ’not’, or ’without’. For example, if a text describes ’a bookshelf with no books, only picture frames’, it should show a bookshelf devoid of books but filled with picture frames instead.
@@
Universal Univ   (Li et al. 2024a) ITI Universal evaluates the ability to accurately depict scenarios where every member of a group shares a specific attribute or is involved in a common relation, as described in a text. It assesses whether the image correctly represents situations indicated by words such as ’every’, ’all’, ’each’, or ’both’. For example, if a text describes ’a bustling kitchen where every chef is preparing a dish’, it should show all chefs within the kitchen actively engaged in cooking.
@@
Complement Comp   (Vempala and Preoţiuc-Pietro 2019) ITI This criterion assesses how text and image work together to enhance or clarify the overall meaning. It evaluates how they complement each other, whether through redundancy, emotional amplification, added context, or interpretation, creating a richer understanding than either could provide alone.
@@
Complementary Overlap ComO   (Vempala and Preoţiuc-Pietro 2019) ITI This criterion captures cases where the image and text redundantly represent the same information, and the image does not contribute anything beyond what is already stated in the text. The text alone is sufficient to convey the core message. For example, a caption saying “Tacos are the best” can be paired with a photo of tacos. Both the image and text communicate the same idea with no additional contextual or semantic contribution from the image.
@@
Emotional Complement EmoC   (Vempala and Preoţiuc-Pietro 2019) ITI This criterion evaluates how the image, particularly through elements like emojis or memes, enhances or clarifies the emotional tone of the text. The image should convey emotions, moods, or subtle feelings that the text alone might not fully express. This could involve using visual elements such as facial expressions, symbols, or humorous images that directly amplify, complement, or interpret the emotional content of the text, helping to create a more vivid or relatable emotional context.
@@
Comment Complement ComC   (Vempala and Preoţiuc-Pietro 2019) ITI This criterion assesses how the text offers commentary or evaluation that adds meaning to the image. The text may provide an interpretation, critique, or personal perspective on the image, offering a deeper understanding or a subjective view of what the image represents. The text can either affirm or challenge the implied meaning of the image, creating a more nuanced interaction between the two modalities.
@@
Contextual Complement ConC   (Vempala and Preoţiuc-Pietro 2019) ITI This criterion evaluates how the text and the image provide complementary background information or contextual details that enhance the understanding of each other. The text should add context, explanations, or relevant background information that helps interpret the image more fully, or conversely, the image should provide visual context that clarifies or enriches the meaning of the text.
Text Repetition TexR (Hu et al. 2024b; Wang et al. 2024b) T Text repetition refers to the unnecessary duplication of phrases, sentences, or ideas that do not add new meaning. Repetition disrupts the flow, reduces clarity, and can make the text feel redundant.
Table 14: The definition and source of Universal Aspects. (Part 5)
@@
Coverage Cov (Li et al. 2023b) Summarization The summary should cover all essential points from the original text, ensuring no significant information is omitted. In this case, completeness is more important than brevity.
Length Constraint LenC (Li et al. 2023b) Summarization This measures the model’s ability to respect any given length restrictions for the response (e.g., sentence count, paragraph count, or specific character limits), ensuring the model does not generate an overly short or excessively long response. If the instruction does not mention a length requirement, all responses should receive a score of -1, without considering the influence of other criteria.
Layout Lay (Li et al. 2023b) Summarization The written summary is encouraged to follow a clear and well-organized layout. You may use headings, bullet points, lists, tables, or other formatting devices to improve readability. This criterion is not applicable if the user does not explicitly require a specific layout.
Conciseness Con (Li et al. 2023b) Summarization This criterion evaluates whether the summary is brief and to the point, effectively capturing the essential information of the original content while avoiding unnecessary details or excessive elaboration. It focuses on the ability to convey key elements clearly and concisely, maintaining the core meaning of the text without including irrelevant or redundant content.
Helpfulness Hel (Li et al. 2024d; Liu et al. 2023a)
@@
Instruction Following InsF (Zeng et al. 2023; Liu et al. 2023a)
@@
Transparency Tran (Zhang et al. 2023b) Medical VQA Transparency is a criterion used to assess how clearly and transparently a model explains its reasoning behind the answer to an academic question. It focuses on the model’s ability to provide understandable, logical, and well-supported justifications for its responses. This includes breaking down complex ideas, referencing relevant information, and demonstrating a step-by-step thought process. A high level of explainability ensures that users not only receive a correct answer but also understand how the answer was derived, helping to build trust in the model’s reasoning and enhancing learning outcomes.
Terminology Term (Zhang et al. 2023b) Medical VQA Terminology evaluates the model’s ability to correctly use and apply medical terminology when responding to queries related to medical image understanding. This includes the accurate usage of terms specific to anatomy, pathology, imaging techniques, or diagnostic processes. For example, if a user asks about a CT scan showing a lung lesion, the model should mention terms like ’pulmonary nodule,’ ’radiographic features,’ or ’lesion characterization’ rather than generic terms like ’spot’ or ’image area.’ Proper use of medical terms indicates the model’s depth of understanding in the field.
Table 15: The definition and source of Task-specific Aspects. (Part 1)
Aspect Abbr Source Sub-Task Definition
Accuracy Acc (Li et al. 2023b) Rewriting Accuracy in Rewriting refers to how correctly the changes follow the given instructions. Any modifications are made according to the task requirements.
Pointing Out PoiO (Li et al. 2023b) Rewriting This criterion evaluates the clarity and transparency of corrections made to the text. Instead of simply returning a corrected version, the correction process must be clearly outlined. This includes pointing out which parts of the text were changed, explaining how the changes were made, and providing a reason for each correction. This approach helps the user understand not only the mistake but also the logic behind the correction, facilitating better learning and comprehension.
Difference Dif (Li et al. 2023b) Rewriting This criterion evaluates whether the written text shows a clear difference from the original in terms of its appearance or structure. The text should not simply be a direct copy of the original; instead, it should reflect meaningful changes, such as rephrasing or reformatting. The goal is to ensure that the new text is distinct and provides added value, rather than just repeating the same content in the same form.
Creativity Cre (Li et al. 2023b)
@@
Theme The (Li et al. 2023b)
@@
Contextual Information ConI (Li et al. 2023b)
@@
Example Quality ExaQ (Li et al. 2023b)
@@
Information Richness InfoR (Li et al. 2023b)
@@
Table 16: The definition and source of Task-specific Aspects. (Part 2)
@@
Suitability Suit (Li et al. 2023b) Title Generation This criterion evaluates how closely the generated title aligns with the content of the text or work. It’s about ensuring that the title directly reflects the main theme or subject matter, providing a meaningful and appropriate summary for the given context. For instance, if the text is about the effects of climate change on wildlife, a relevant title might be Wildlife in Distress: Unraveling the Impact of Climate Change.
Coverage Cov (Li et al. 2023b) Title Generation This criterion evaluates how much of the essential information from the source content is captured in the generated title. It’s about assessing how well the title encapsulates the key points or main ideas of the source content. For instance, if the text discusses the history, types, and health benefits of meditation, a title that covers these points might be Meditation: Tracing Its Roots and Understanding Its Health Wonders.
Appeal Appe (Li et al. 2023b) Title Generation This criterion evaluates the attractiveness of the generated title. It assesses whether the title has the power to catch the reader’s attention, spark interest, and entice the reader to delve into the content. For example, a title like ’Unlocking Creativity: Harnessing the Power of an Unconventional Mind’ might appeal to readers interested in personal development and creativity.
Translation Fidelity TrnF (Li et al. 2023b) Translation This criterion evaluates the precision of the translation in replicating the source text’s content in the target language. It involves correct vocabulary choices, proper grammar, and syntax. It ensures that the meaning, tone, and nuances of the original text are preserved without adding, omitting, or distorting information.
Thoroughness Thor (Li et al. 2023b) Translation This criterion assesses whether the translation covers all parts of the source text without leaving out any information, details, or nuances. It ensures that the translation fully conveys the original text’s content, including all its ideas, arguments, and points, in the target language.
Precision Prec (Li et al. 2023b) Data Analysis This criterion evaluates the precision and correctness of the data analysis results. It assesses the degree to which the results correspond to the true or actual values, ensuring that the output is free from errors or distortions.
Format Form (Li et al. 2023b) Data Analysis This criterion evaluates the clarity and user-friendliness of the data analysis output generated by the model. It assesses whether the output is well-structured, follows standard data presentation conventions, and is easy to interpret and understand. The output should present the analysis results in a clear and concise manner, using appropriate tables, charts, or other visual aids as necessary. The use of clear headings, labels, and explanations should also be considered.
@@
Engagingness Eng (Li et al. 2023b) Question Generation This criterion evaluates the ability of the generated questions to captivate and hold the attention of the audience. Engaging questions spark curiosity, stimulate thought, and encourage active participation. For instance, a question like ’How does climate change impact our daily lives?’ might engage readers because it connects a global issue to personal experiences.
@@
Difficulty Diff (Li et al. 2023b) Question Generation This criterion assesses whether the complexity level of the generated questions matches the knowledge and comprehension level of the intended audience. For example, for a high school audience, a question like ’Can you explain Einstein’s theory of relativity?’ might be too difficult, while ’What is the capital of France?’ might be too easy.
@@
Relevance Rel (Li et al. 2023b) Keywords Extraction In the context of keyword extraction, this criterion evaluates how closely the extracted keywords align with the main themes or subjects of the text. Relevant keywords should accurately reflect the core ideas, topics, or concepts presented in the text, providing a concise and meaningful representation of the content.
@@
Representation Repr (Li et al. 2023b) Keywords Extraction In the context of keyword extraction, this criterion assesses the extent to which the extracted keywords represent all the significant themes, topics, or concepts in the text. A high coverage means that the keywords collectively provide a comprehensive summary of the text, leaving no important aspect unrepresented.
Table 17: The definition and source of Task-specific Aspects. (Part 3)
Aspect Abbr Source Sub-Task Definition
Accuracy Acc (Li et al. 2024d; Zhang et al. 2023c)
@@
Instruction Following InsF (Li et al. 2024d; Zeng et al. 2023)
@@
Instruction Inconsistency InsI (Li et al. 2024d; Huang et al. 2025)
@@
Context Inconsistency ConI (Li et al. 2024d; Huang et al. 2025)
@@
Instruction Following InsF (Zeng et al. 2023; Masry et al. 2022)
@@
Instruction Following InsF (Zeng et al. 2023; Mathew et al. 2022)
@@
Instruction Following InsF (Zeng et al. 2023; Singh et al. 2019)
@@
Table 18: The definition and source of Task-specific Aspects. (Part 4)
Aspect Abbr Source Sub-Task Definition
Conciseness Con (Mokady, Hertz, and Bermano 2021; Hossain et al. 2019)
@@
Object Alignment ObjA (Mokady, Hertz, and Bermano 2021; Hossain et al. 2019)
@@
Text-Image Relationship T-IR (Mokady, Hertz, and Bermano 2021; Hossain et al. 2019)
@@
Alignment Ali (Mokady, Hertz, and Bermano 2021; Hossain et al. 2019)
@@
Completeness Com (Hossain et al. 2019; Zhao et al. 2023)
@@
Attribute Alignment AttA (Hossain et al. 2019; Zhao et al. 2023)
@@
Spatial Alignment SpaA (Hossain et al. 2019; Zhao et al. 2023)
@@
Count Alignment CouA (Hossain et al. 2019; Zhao et al. 2023)
@@
Action Alignment ActA (Hossain et al. 2019; Zhao et al. 2023)
@@
Object Alignment ObjA (Hossain et al. 2019; Zhao et al. 2023)
@@
Table 19: The definition and source of Task-specific Aspects. (Part 5)
@@
Color Alignment ColA (Hossain et al. 2019; Zhao et al. 2023)
@@
Texture Alignment TexA (Hossain et al. 2019; Zhao et al. 2023)
@@
Shape Alignment ShaA (Hossain et al. 2019; Zhao et al. 2023)

### v4

@@ C.3 Aspects Used in the public benchmark
Shape Alignment ShaA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023) ITI This criterion evaluates how accurately the shapes described in the text correspond to the shapes of objects in the image, based solely on the text’s shape descriptors, or whether the shape of the objects conforms to realistic shape. The text may include common shape terms such as square, round, rectangular, oval, triangular, cubic, cylindrical, spherical, and other basic or specialized shapes like pyramidal, conical, pentagonal, teardrop, crescent, and diamond.
@@
Texture Alignment TexA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023) ITI This criterion evaluates how accurately the textures described in the text are reflected in the image, based solely on the text description, or whether the textures of the objects conform to realistic textures. It includes surface qualities such as smoothness, roughness, or granularity, as well as material types like wood, plastic, and rubber. Additionally, it may encompass tactile sensations or patterns, such as silky, polished, or other material-based descriptions. The accuracy of the texture description is assessed by comparing the textual description to the actual appearance and surface properties of the objects in the image.
@@
Advanced Alignment AdvA (Li et al., 2024a) ITI Advanced Alignment is a comprehensive evaluation aspect that assesses the ability to accurately represent complex textual descriptions in an image. It encapsulates Comparison, Negation, and Universality aspects. This means it checks whether the image correctly depicts comparisons between entities based on attributes such as number or volume; represents absence or contradiction of elements indicated by words like ’no’ or ’not’; and illustrates scenarios where every member of a group shares a specific attribute or relation.
@@
Comparison Cmpa (Li et al., 2024a) ITI Comparison evaluates the ability to accurately depict comparisons between entities as described in a text. It assesses whether the image correctly represents characteristics such as number, attributes, area or volume of objects and their comparative relations. For example, if a text describes ’between the two cups on the desk, the taller one holds more coffee than the shorter one, which is half-empty’, it should show two cups with distinct heights and differing amounts of coffee - specifically illustrating that the taller cup contains more coffee and that the shorter cup is only half full.
@@
Negation Nega (Li et al., 2024a) ITI Negation evaluates the ability to accurately depict the absence or contradiction of elements as described in a text. It assesses whether the image correctly represents scenarios where entities are not present, as indicated by words such as ’no’, ’not’, or ’without’. For example, if a text describes ’a bookshelf with no books, only picture frames’, it should show a bookshelf devoid of books but filled with picture frames instead.
@@
Universal Univ   (Li et al., 2024a) ITI Universal evaluates the ability to accurately depict scenarios where every member of a group shares a specific attribute or is involved in a common relation, as described in a text. It assesses whether the image correctly represents situations indicated by words such as ’every’, ’all’, ’each’, or ’both’. For example, if a text describes ’a bustling kitchen where every chef is preparing a dish’, it should show all chefs within the kitchen actively engaged in cooking.
@@
Complement Comp   (Vempala & Preoţiuc-Pietro, 2019) ITI This criterion assesses how text and image work together to enhance or clarify the overall meaning. It evaluates how they complement each other, whether through redundancy, emotional amplification, added context, or interpretation, creating a richer understanding than either could provide alone.
@@
Complementary Overlap ComO   (Vempala & Preoţiuc-Pietro, 2019) ITI This criterion captures cases where the image and text redundantly represent the same information, and the image does not contribute anything beyond what is already stated in the text. The text alone is sufficient to convey the core message. For example, a caption saying “Tacos are the best” can be paired with a photo of tacos. Both the image and text communicate the same idea with no additional contextual or semantic contribution from the image.
@@
Emotional Complement EmoC   (Vempala & Preoţiuc-Pietro, 2019) ITI This criterion evaluates how the image, particularly through elements like emojis or memes, enhances or clarifies the emotional tone of the text. The image should convey emotions, moods, or subtle feelings that the text alone might not fully express. This could involve using visual elements such as facial expressions, symbols, or humorous images that directly amplify, complement, or interpret the emotional content of the text, helping to create a more vivid or relatable emotional context.
@@
Comment Complement ComC   (Vempala & Preoţiuc-Pietro, 2019) ITI This criterion assesses how the text offers commentary or evaluation that adds meaning to the image. The text may provide an interpretation, critique, or personal perspective on the image, offering a deeper understanding or a subjective view of what the image represents. The text can either affirm or challenge the implied meaning of the image, creating a more nuanced interaction between the two modalities.
@@
Contextual Complement ConC   (Vempala & Preoţiuc-Pietro, 2019) ITI This criterion evaluates how the text and the image provide complementary background information or contextual details that enhance the understanding of each other. The text should add context, explanations, or relevant background information that helps interpret the image more fully, or conversely, the image should provide visual context that clarifies or enriches the meaning of the text.
Table 16: The definition and source of Task-specific Aspects (TAs). (Part 1)
@@
Coverage Cov (Li et al., 2023b) Summarization The summary should cover all essential points from the original text, ensuring no significant information is omitted. In this case, completeness is more important than brevity.
Length Constraint LenC (Li et al., 2023b) Summarization This measures the model’s ability to respect any given length restrictions for the response (e.g., sentence count, paragraph count, or specific character limits), ensuring the model does not generate an overly short or excessively long response. If the instruction does not mention a length requirement, all responses should receive a score of -1, without considering the influence of other criteria.
Layout Lay (Li et al., 2023b) Summarization The written summary is encouraged to follow a clear and well-organized layout. You may use headings, bullet points, lists, tables, or other formatting devices to improve readability. This criterion is not applicable if the user does not explicitly require a specific layout.
Conciseness Con (Li et al., 2023b) Summarization This criterion evaluates whether the summary is brief and to the point, effectively capturing the essential information of the original content while avoiding unnecessary details or excessive elaboration. It focuses on the ability to convey key elements clearly and concisely, maintaining the core meaning of the text without including irrelevant or redundant content.
Helpfulness Hel (Li et al., 2024f; Liu et al., 2023a)
@@
Instruction Following InsF (Zeng et al., 2023; Liu et al., 2023a)
@@
Transparency Tran (Zhang et al., 2023b) Medical VQA Transparency is a criterion used to assess how clearly and transparently a model explains its reasoning behind the answer to an academic question. It focuses on the model’s ability to provide understandable, logical, and well-supported justifications for its responses. This includes breaking down complex ideas, referencing relevant information, and demonstrating a step-by-step thought process. A high level of explainability ensures that users not only receive a correct answer but also understand how the answer was derived, helping to build trust in the model’s reasoning and enhancing learning outcomes.
Terminology Term (Zhang et al., 2023b) Medical VQA Terminology evaluates the model’s ability to correctly use and apply medical terminology when responding to queries related to medical image understanding. This includes the accurate usage of terms specific to anatomy, pathology, imaging techniques, or diagnostic processes. For example, if a user asks about a CT scan showing a lung lesion, the model should mention terms like ’pulmonary nodule,’ ’radiographic features,’ or ’lesion characterization’ rather than generic terms like ’spot’ or ’image area.’ Proper use of medical terms indicates the model’s depth of understanding in the field.
Accuracy Acc (Li et al., 2023b) Rewriting Accuracy in Rewriting refers to how correctly the changes follow the given instructions. Any modifications are made according to the task requirements.
Pointing Out PoiO (Li et al., 2023b) Rewriting This criterion evaluates the clarity and transparency of corrections made to the text. Instead of simply returning a corrected version, the correction process must be clearly outlined. This includes pointing out which parts of the text were changed, explaining how the changes were made, and providing a reason for each correction. This approach helps the user understand not only the mistake but also the logic behind the correction, facilitating better learning and comprehension.
Difference Dif (Li et al., 2023b) Rewriting This criterion evaluates whether the written text shows a clear difference from the original in terms of its appearance or structure. The text should not simply be a direct copy of the original; instead, it should reflect meaningful changes, such as rephrasing or reformatting. The goal is to ensure that the new text is distinct and provides added value, rather than just repeating the same content in the same form.
Creativity Cre (Li et al., 2023b)
@@
Table 17: The definition and source of Task-specific Aspects (TAs). (Part 2)
Aspect Abbr Source Sub-Task Definition
Theme The (Li et al., 2023b)
@@
Contextual Information ConI (Li et al., 2023b)
@@
Example Quality ExaQ (Li et al., 2023b)
@@
Information Richness InfoR (Li et al., 2023b)
@@
Suitability Suit (Li et al., 2023b) Title Generation This criterion evaluates how closely the generated title aligns with the content of the text or work. It’s about ensuring that the title directly reflects the main theme or subject matter, providing a meaningful and appropriate summary for the given context. For instance, if the text is about the effects of climate change on wildlife, a relevant title might be Wildlife in Distress: Unraveling the Impact of Climate Change.
Coverage Cov (Li et al., 2023b) Title Generation This criterion evaluates how much of the essential information from the source content is captured in the generated title. It’s about assessing how well the title encapsulates the key points or main ideas of the source content. For instance, if the text discusses the history, types, and health benefits of meditation, a title that covers these points might be Meditation: Tracing Its Roots and Understanding Its Health Wonders.
Appeal Appe (Li et al., 2023b) Title Generation This criterion evaluates the attractiveness of the generated title. It assesses whether the title has the power to catch the reader’s attention, spark interest, and entice the reader to delve into the content. For example, a title like ’Unlocking Creativity: Harnessing the Power of an Unconventional Mind’ might appeal to readers interested in personal development and creativity.
Translation Fidelity TrnF (Li et al., 2023b) Translation This criterion evaluates the precision of the translation in replicating the source text’s content in the target language. It involves correct vocabulary choices, proper grammar, and syntax. It ensures that the meaning, tone, and nuances of the original text are preserved without adding, omitting, or distorting information.
Thoroughness Thor (Li et al., 2023b) Translation This criterion assesses whether the translation covers all parts of the source text without leaving out any information, details, or nuances. It ensures that the translation fully conveys the original text’s content, including all its ideas, arguments, and points, in the target language.
Precision Prec Li et al. (2023b) Data Analysis This criterion evaluates the precision and correctness of the data analysis results. It assesses the degree to which the results correspond to the true or actual values, ensuring that the output is free from errors or distortions.
Format Form (Li et al., 2023b) Data Analysis This criterion evaluates the clarity and user-friendliness of the data analysis output generated by the model. It assesses whether the output is well-structured, follows standard data presentation conventions, and is easy to interpret and understand. The output should present the analysis results in a clear and concise manner, using appropriate tables, charts, or other visual aids as necessary. The use of clear headings, labels, and explanations should also be considered.
Table 18: The definition and source of Task-specific Aspects (TAs). (Part 3)
@@
Engagingness Eng (Li et al., 2023b) Question Generation This criterion evaluates the ability of the generated questions to captivate and hold the attention of the audience. Engaging questions spark curiosity, stimulate thought, and encourage active participation. For instance, a question like ’How does climate change impact our daily lives?’ might engage readers because it connects a global issue to personal experiences.
@@
Difficulty Diff (Li et al., 2023b) Question Generation This criterion assesses whether the complexity level of the generated questions matches the knowledge and comprehension level of the intended audience. For example, for a high school audience, a question like ’Can you explain Einstein’s theory of relativity?’ might be too difficult, while ’What is the capital of France?’ might be too easy.
@@
Relevance Rel (Li et al., 2023b) Keywords Extraction In the context of keyword extraction, this criterion evaluates how closely the extracted keywords align with the main themes or subjects of the text. Relevant keywords should accurately reflect the core ideas, topics, or concepts presented in the text, providing a concise and meaningful representation of the content.
@@
Representation Repr (Li et al., 2023b) Keywords Extraction In the context of keyword extraction, this criterion assesses the extent to which the extracted keywords represent all the significant themes, topics, or concepts in the text. A high coverage means that the keywords collectively provide a comprehensive summary of the text, leaving no important aspect unrepresented.
Accuracy Acc (Li et al., 2024f; Zhang et al., 2023c)
@@
Instruction Following InsF (Li et al., 2024f; Zeng et al., 2023)
@@
Instruction Inconsistency InsI (Li et al., 2024f; Huang et al., 2025)
@@
Context Inconsistency ConI (Li et al., 2024f; Huang et al., 2025)
@@
Instruction Following InsF (Zeng et al., 2023; Masry et al., 2022)
@@
Instruction Following InsF (Zeng et al., 2023; Mathew et al., 2022)
@@
Reasonableness Reas (Liu et al., 2024a)
Storytelling
Generation
Table 19: The definition and source of Task-specific Aspects (TAs). (Part 4)
Aspect Abbr Source Sub-Task Definition
Instruction Following InsF (Zeng et al., 2023; Singh et al., 2019)
@@
Conciseness Con (Mokady et al., 2021; Hossain et al., 2019)
@@
Object Alignment ObjA (Mokady et al., 2021; Hossain et al., 2019)
@@
Text-Image Relationship T-IR (Mokady et al., 2021; Hossain et al., 2019)
@@
Alignment Ali (Mokady et al., 2021; Hossain et al., 2019)
@@
Completeness Com (Hossain et al., 2019; Zhao et al., 2023)
@@
Attribute Alignment AttA (Hossain et al., 2019; Zhao et al., 2023)
@@
Spatial Alignment SpaA (Hossain et al., 2019; Zhao et al., 2023)
@@
Count Alignment CouA (Hossain et al., 2019; Zhao et al., 2023)
@@
Action Alignment ActA (Hossain et al., 2019; Zhao et al., 2023)
@@
Object Alignment ObjA (Hossain et al., 2019; Zhao et al., 2023)
@@
Table 20: The definition and source of Task-specific Aspects (TAs). (Part 5)
@@
Color Alignment ColA (Hossain et al., 2019; Zhao et al., 2023)
@@
Texture Alignment TexA (Hossain et al., 2019; Zhao et al., 2023)
@@
Shape Alignment ShaA (Hossain et al., 2019; Zhao et al., 2023)